# Construction of Functional Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

### v1

@@ Abstract
The convergence of materials science and artificial intelligence has unlocked new opportunities for gathering, analyzing, and generating novel materials sourced from extensive scientific literature. Despite the potential benefits, persistent challenges such as manual annotation, precise extraction, and traceability issues remain. Large language models have emerged as promising solutions to address these obstacles. This paper introduces Functional Materials Knowledge Graph (FMKG), a multidisciplinary materials science knowledge graph. Through the utilization of advanced natural language processing techniques, extracting millions of entities to form triples from a corpus comprising all high-quality research papers published in the last decade. It organizes unstructured information into nine distinct labels, covering Name, Formula, Acronym, Structure/Phase, Properties, Descriptor, Synthesis, Characterization Method, Application, and Domain, seamlessly integrating papers’ Digital Object Identifiers. As the latest structured database for functional materials, FMKG acts as a powerful catalyst for expediting the development of functional materials and a fundation for building a more comprehensive material knowledge graph using full paper text. Furthermore, our research lays the groundwork for practical text-mining-based knowledge management systems, not only in intricate materials systems but also applicable to other specialized domains.
@@ Introduction
Furthermore, the complexity of materials information extends beyond composition and structure to encompass their respective application fields. For instance, organic materials are commonly utilized in biological applications, semiconductors serve as integral components in electronics, and metals find applications in mechanical engineering. The process of designing novel materials typically commences with a clear understanding of their intended applications, aiming to maximize research efficiency. Given this context, compared to universal databases like the Materials Project, which might not be as useful due to their broad focus, specific databases focused on applications or traits hold the potential to provide more valuable information for researchers in relevant industries. This underscores the importance of developing dedicated databases that cater to the nuanced needs of materials science research, facilitating a more targeted approach to material discovery and application.
@@
In this paper, we have achieved significant advancements in the development of Functional Materials Knowledge Graph (FMKG), a pioneering graph database tailored for the field of functional materials. Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first knowledge graph dedicated for functional materials, where researchers can easily get the information about functional material through query the FMKG. 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs.
@@ Methods
Figure 1(a) illustrates the comprehensive workflow of our research. Through NER and RE tasks, we extract the structural information about catalyst, battery and solar cell. After ER and normalization, we integrate information from these three fields and construct a knowledge graph. Specifically, the pipeline displayed in the Figure 1(b), commencing with the manual annotation and normalization of the initial training data set, prompting the fine-tuned LLMs specifically for NER and RE tasks. This inference dataset is subsequently divided into ten batches, a crucial step for the iterative process that follows. Then, we finish ER task through the NLP technology including ChemDataExactor [22], mat2vec [23] and our expert dictionary. After ER, high-quality results are meticulously selected to augment the training set, thereby enhancing the model’s performance in subsequent iterations. Finally, the knowledge graph is constructed using the triples transferred from normalizd result after the last iteration.
@@
Material experts undertook the annotation of nine distinct categories from 75 abstracts of research papers, act as LLM’s training dataset. These categories encompassed core labels - "Name," "Formula," or "Acronym" - which consistently represent specific materials across diverse papers, albeit in varied forms. Additionally, supplementary labels included "Descriptor," "Structure/Phase," "Application," "Property," "Synthesis," and "Characterization." The "Property" label served to delineate quantitative properties such as ’length,’ ’specific surface area,’ and ’mass,’ whereas the "Descriptor" label encapsulated qualitative attributes like ’stable,’ ’vertically,’ and ’safe.’ Since from the application aspects, developers always concern about its own properties, we also annotated the essential parameters as "Property" directly linked to specific "Application". For instance, attributes like ’specific capacity’ for lithium-ion batteries, ’energy conversion efficiency’ for solar cells, and ’H2 production rate’ for hydrogen evolution reaction were systematically annotated.
As the Figure 2 shows, the central node, "Material," is pivotal and connects to nodes that describe its nomenclature, composition, and various attributes. Specifically, "Material" is consisted of "Formula", "Name" and "Acronym". "Structure/Phase" describing its physical form; "Application" denoting its practical uses; and "Property" outlining its inherent characteristics. The "Material" node is also associated with "Descriptor" node that provide additional qualitative information. Moreover, the "Application" node branches out to "Property", "Descriptor" and a "Domain" suggesting further specification and contextual relevance of application. To save the source of nodes and relations, each node is connected to a "Digital Object Identifier (DOI)" node. If we want to determine the source article from which the relation is derived, we can obtain the result by querying the intersection of "DOI" node neighbors at both ends of the relation.
Figure 2: This schematic represents the FMKG schema (a) All the nodes in FMKG connect to all their sources - "DOI". (b) Relations between all labels. (c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured uniform representation of entities with similar meanings. For example, terms such as ’Lithium-Ion Battery’ and ’Li-ion batteries’ were standardized to ’lithium-ion battery,’ while phrases like ’solution casting method,’ ’solvent post-treatment method,’ and ’solution-based deposition’ were simplified to ’solution-processed.’ All annotated entities underwent this normalization process, ensuring consistency and facilitating effective training of LLMs. The field of functional materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
@@
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA7b, LLaMA2 7b, and Darwin [24]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
Precision=True Positives (tp)True Positives (tp) + False Positives ($\text{Precision}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) +False Positives (fp)}}$ (1)
Recall=True Positives (tp)True Positives (tp) + False Negatives ($\text{Recall}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) + False Negatives (fn)}}$ (2)
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
For NER, a true positive (tp) is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema. Conversely, a false positive (fp) occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative (fn) is when the model fails to identify a correct entity, relation, or schema element that should have been recognized. After definition, we can evaluate each task using Equations (1), (2), and (3).
@@
The quality of KG is crucial for its credibility, we need to further check and correct the inference results before graph construction. For the core label, we divide standardization into two steps: ER-NF/A to ER-N/F. In stage of ER-NF/A, we apply the ChemDataExtractor to derive the Formula and pairs of "Name - Acronym" from the abstract. Subsequently, we utilize the mat2vec model to perform word embedding on entities extracted by both the LLM and ChemDataExtractor. Through comparing their similarities, we aim to correct the core entities to guarantee their precision and the accurate correspondence between "Name" and "Acronym". However, The Name and Formula of materials are often placed in incorrect labels and are difficult to distinguish through ChemDataExtractor. Therefore, we conduct ER-N/F. We selected 2000 correctly classified labels as the training set and used to fine-tune LLM to complete a binary classification problem. Additionally, we selected other 200 labels to evaluate the performance of binary classification.
To normalize entities and relations in other label (ER-OL), we conduct a Word Embedding Clustering to create a expert dictionary. We convert words into vector form using mat2vec and group semantically similar words into the same cluster based on their vector similarity. Specifically, we develop a Density-Based Dynamic Vector Clustering method that dynamically forms clusters based on the similarity of vector representations, without predefining the number of clusters. This approach utilizes a similarity threshold to determine cluster membership, allowing entities to join existing clusters or form new ones based on their proximity in the feature space, effectively adapting to the inherent structure of the data.
By naming each cluster through experts in the field of material science, we extracted a total of approximately 600 vocabulary, including the structure, phase, applications, characterization methods, synthesis methods, properties, and descriptions of most energy materials. Subsequently, we adopt analytic strategies and embed the vocabulary from expert dictionary and the enetities extracted by LLM using mat2vec. After similarity comparison, we finish the standardization and ensure that both entities and relations are correct. It is worth mentioning that the content of "Property" and "Descriptor" is often not limited to the energy field and may cover more information. Therefore, we implement relatively tolerant standardization for these two labels, which only normalize them with matching content but does not delete unmatched entities and relations.
To improve the caliber of the training dataset, we select a small number of high-quality data from the normalized inference output of each iterative process, incorporating this data into the training set. High-quality data is characterized by its high precision and recall metrics. Furthermore, the performance of the fine-tuned LLM, trained on each progressively augmented training set, is rigorously evaluated to ensure optimal effectiveness.
@@
To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head, the names of other labels forming the relations, and their values acting as tails. Within the hierarchy of core labels, "Formula" is accorded the highest priority, followed by "Name", and then "Acronym". Furthermore, each head and tail node is linked to the DOI of the source article associated with the triplet. This setup allows us to ascertain the provenance of the relation between any two nodes by examining the intersection in their connected entities. Then we transfer the triples into FMKG and store the FMKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions. To facilitate access to the detailed information, we also make the dataset available in the CSV format for straightforward data handling.
@@
In Table 1, the performance comparison of LLMs includes Darwin, LLaMA, and LLaMA2 in NER, RE and ER task within the materials science domain is presented. Darwin achieves markedly higher F1 scores in both NER and RE tasks than the LLaMA 7b and LLaMA2 7b models. This indicates that Darwin can perform more satisfactory results in text tasks in the field of materials science. However, there has been no significant difference in their effectiveness for ER tasks between these models. We speculate that this is due to insufficient contextual memory capabilities of the LLMs.
In response to the challenge, entity resolution was applied to the inference outputs, which substantially improved the performance in Standardization (ER) tasks. This refinement process has also led to enhancements in the NER and Relation RE tasks as well. The improvements can be attributed to the ER task removing some of the incorrect inferences that resulted from hallucinations and providing a secondary adjustment of the extracted relations.
@@
Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
To gain a more comprehensive understanding of the improvements brought by each step in the standardization process, we conducted ablation experiments, the results are shown in Table 2 which clearly shows that each method contributes to the improvement of ER. The result indicates the most effective method is to optimize through the expert dictionary, which indicates that in addition to the accuracy of the expert dictionary, it can also fully cover all other labels except for the three labels representing the material. The improvement of ER-NF/A is also significant, as this process can eliminate most erroneous materials, and even has a stronger improvement on NER than ER-OL.
@@
ER-OL 88.59 -4.37 80.20 -8.07 50.00 -38.83
Table 2: Result of the ablation experiment in normalization.
Besides, ER-N/F pay attention to adjust the relation error between "Name" and "Formula", from a contribution perspective, it is also an indispensable part of ER. We also evaluated the LLM classifier by selecting 200 manually labeled data, and the statistical evaluation results in Table 3 shows the strong performance of LLM used in classification task.
Model Accuracy F1 Score Specificity
Darwin 0.9650 0.9581 0.9741
Table 3: Binary classification performance through Fine-tuned Darwin.
The triples are transferred from normalized data to construct the FMKG which contains 162,605 nodes and 731,772 edges, the schematic diagram shows in Figure 3. In the graph, there are 11 types of nodes and 13 types of relations. The FMKG store large amount of information in functional materials includes battery, catalyst and solar cell. For example, as the Figure 4 shows, by analyzing the frequency of material appearances in the battery domain, we observe that Co2O3 is the most frequently occurring material, followed by MoS2, graphite, TiO2, LiCoO2 etc which are commonly used in the battery domain. Likewise, it is evident that the most prevalent application within the battery field is lithium-ion batteries, which appear with a frequency significantly higher than other applications in the battery domain. These insights align with our understanding of these domains and materials, confirming that our knowledge graph effectively stores factual knowledge and offers a knowledge platform for material researchers.
Figure 3: (a) Global schematic diagram of FMKG; (b) Local schematic diagram of FMKG.
Figure 4: The image shows the top ten materials and applications in each field.
For further validation of FMKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, annotator can obtain 1000 entities and 500 relations. The report from annotator shows in Table 4.
“Name” is the highest priority core label and only serves as the head in the knowledge graph, so it will not be used as a relation. Notably, the labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous standardization of entities under these categories, using our expert dictionary. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation. But the results are still satisfactory.
"Name" and "Acronym" has also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. The impact of this kind of error is not significant, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.

### v2

# Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

@@ Abstract
Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges for efficient discovery and integration of new materials. Traditional methods, often reliant on costly and time-consuming experimental approaches, further complicate rapid innovation. Addressing these challenges, the integration of artificial intelligence with materials science has opened avenues for accelerating the discovery process, though it also demands precise annotation, data extraction, and traceability of information. To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques, integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges. MKG categorizes information into comprehensive labels such as Name, Formula, and Application, structured around a meticulously designed ontology, thus enhancing data usability and integration. By implementing network-based algorithms, MKG not only facilitates efficient link prediction but also significantly reduces reliance on traditional experimental methods. This structured approach not only streamlines materials research but also lays the groundwork for more sophisticated science knowledge graphs.
@@ Introduction
In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science. Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first accurate knowledge graph dedicated for materials, where researchers can easily get the information about material through query the MKG. 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs. 4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
@@ Methods
Figure 1(a) presents the elaborate workflow of our study. Initially, through Named Entity Recognition (NER) and Relation Extraction (RE) tasks, we extract structural information pertinent to catalysts, batteries, and solar cells. Subsequent Entity Resolution (ER) and data normalization steps enable the integration of information from these distinct fields into a unified knowledge graph. Specifically, as outlined in Figure 1(b), the process begins with the manual annotation and normalization of the initial training dataset, essential for fine-tuning Large Language Models (LLMs) tailored for NERRE tasks. The inference dataset is then methodically divided into ten batches, facilitating the iterative refinement process that follows. The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [22], mat2vec [23], and our expertly curated dictionary. After ER, we selectively enhance the training dataset with high-quality results to improve the model’s performance in subsequent iterations. The knowledge graph is finally constructed using the normalized data from the last iteration. To complete the graph and predict potential material applications, we employ both network-based algorithms and graph embeddings. This methodology provides critical insights and recommendations for researchers in the materials science domain.
@@
Material experts annotated nine distinct categories from the abstracts of 75 research papers, forming the training dataset for the Large Language Models (LLMs). The structure of the data reflects the structure of the KG. As illustrated in Figure 2, the "Material" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes. Specifically, the "Material" node comprises sub-nodes such as "Formula," "Name," and "Acronym" for identifying materials. The "Structure/Phase" node describes the material’s physical state, while the "Application" node denotes its practical uses, and the "Property" node details its inherent characteristics. Additionally, the "Descriptor" node provides qualitative information enhancing the contextual understanding of the material.
Furthermore, the "Application" node is extended to include "Property", "Descriptor", and "Domain" nodes, indicating the specific attributes and the broader context of the material’s application. To maintain data provenance and traceability, each node is linked to a "Digital Object Identifier (DOI)" node. This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations was derived, ensuring data integrity within the graph structure.
Figure 2: This schema represents the ontology of MKG (a) All the nodes in MKG connect to all their sources - "DOI". (b) Relations between all labels. (c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured uniform representation of entities with similar meanings. For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed", ensuring consistency and facilitating effective training of LLMs. The field of materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
@@
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [24], and Darwin [25]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
For NER, a true positive is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema. Conversely, a false positive occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative is when the model fails to identify a correct entity, relation, or schema element that should have been recognized. After defining these terms, we can evaluate each task using standard precision, recall, and F1 score metrics.
@@
The quality of KG is crucial for its credibility, to further check and correct the inference results before graph construction. To ensure the precision of these results, we initially employ ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts. Subsequently, entities recognized by both the Large Language Model (LLM) and ChemDataExtractor are embedded using the mat2vec model. We analyze their similarities to rectify core entities and ensure accurate "Name" to "Acronym" associations. Given the frequent mislabeling of "Name" and "Formula", we progress to the Entity Resolution - Name/Formulas (ER-N/F) phase. Here, we refine the LLM using a specifically curated training set comprising 2,000 accurately labeled entities for binary classification, which is evaluated against an additional set of 200 labels.
For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [26] algorithm to create the expert dictionary (ER-ED). This algorithm dynamically forms clusters based on vector similarity, without the need to predefine the number of clusters. Each cluster is named by material science experts, leading to the creation of an expert dictionary containing approximately 600 terms related to structures, phases, applications, and more. This dictionary, along with the entities extracted by the LLM, undergoes similarity analysis to standardize and confirm the accuracy of both entities and relations. To elevate the quality of the training dataset, we selectively integrate high-quality data from normalized inference outputs from each iteration into the training set, continuously monitoring and enhancing the performance of the fine-tuned LLM to ensure its efficacy.
@@
Given a collection of entities $E$ and relations $R$, a knowledge graph $\mathcal{K}=E\times R\times E$ is structured as a directed multi-relational graph. It comprises triples formatted as $(h,r,t)\in\mathcal{K}$, where $h$ and $t$ are entities within $E$. To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head $h$, the names of other labels forming the relations $R$, and their values acting as tails $t$. Within the hierarchy of core labels, "Formula" is accorded the highest priority, followed by "Name", and then "Acronym". Furthermore, each head $h$ and tail $t$ node is linked to the DOI of the source article associated with the triplet. This setup allows us to ascertain the provenance of the relation between any two nodes by examining the intersection in their connected entities. Then we transfer the triples into MKG and store the MKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions. To facilitate access to the detailed information, we also make the dataset available in the RDF and CSV format for straightforward data handling.
Graph completion
In the analysis depicted in Figure 3, we enhance the traditional Jaccard similarity, hereafter referred to as $J$, to form a comprehensive similarity metric for evaluating potential material-to-application links. The Jaccard similarity is mathematically defined as:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
where $A$ and $B$ represent the attribute sets of two compared materials, such as their properties or applications. To refine our predication method, we extend this basic formulation by incorporating weighted factors that reflect the relevance and impact of each attribute, resulting in a modified similarity score for "Materials-Applications", $J_{\text{mat}}$. This score is computed as follows:
$J_{\text{mat}}(A,B)=\sum_{i=1}^{n}w_{i}\cdot\frac{|A_{i}\cap B_{i}|}{|A_{i}\cup B_{i}|}$
where $n$ is the number of attributes considered, $A_{i}$ and $B_{i}$ are the specific attributes of materials $A$ and $B$ respectively, and $w_{i}$ are the weights assigned to each attribute, signifying their importance in the context of material application. This approach mimics the investigative processes commonly employed by materials scientists when developing new materials. In determining the suitability of new materials for specific applications, researchers systematically analyze their chemical and physical properties, as well as their structural characteristics, comparing these attributes to those of materials well-established in the target field. This enhanced method not only accounts for the quantity of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application. A higher $J_{\text{modified}}$ score indicates a stronger likelihood of applicability between the material in question and the target application. Thus, materials scoring high on this metric are considered promising candidates for the specified applications.
Figure 3: The process of network-based graph completion.
In addition to the Enhanced Jaccard Similarity, we employ the TransE model [27]. TransE treats entities and relations as vectors in the same embedding space. The core idea of TransE is to model relations by interpreting them as translations in the embedding space. For a given relation $r$ between head $h$ and tail $t$, the relationship is said to hold if:
$\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|\approx 0$
[…]
@@
Figure 4: Schematic comparison of MKG and MatKG2.
To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [15], highlighting the differences in their construction methodologies as illustrated in Figure 4. MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation. In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously. This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in the Table 1. Darwin significantly outperforms both the LLaMA 7b and LLaMA2 7b models in the F1 scores for NER and RE tasks, suggesting that Darwin yields more effective results in text-related tasks in materials science. However, there is no marked difference in the performance of these models on ER tasks, we can say LLM has a weaker ability to complete ER in our task. This may be attributed to the limited contextual memory capabilities of the LLMs and this is we need to apply additional ER process. The normailzed Darwin shows the performance of our ER and normalization process, the result indicate that it not only achieve the ER task successfully, but also contribute to the NER and RE task.
Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
@@
To better understand the enhancements provided by each component in the ER/Normalization process, ablation studies were performed. The outcomes, detailed in Table 2, demonstrate that each technique positively affects the effectiveness of ER. The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels which are included in the dictionary. Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExactor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED. Furthermore, from the results of the ablation experiment, ER-N/F appears that its contribution is not as significant as the other two components, but considering the confusion between the material Formula and Name may have an impact on subsequent MKG applications. ER-N/F is also an indispensable part of the ER process. The normalized data is transformed into triples to construct the MKG, which comprises 162,605 nodes and 731,772 edges, as illustrated in the schematic diagram in Figure 5.
Table 2: Result of the ablation experiment in normalization.
@@
ER-ED 88.59 -4.37 80.20 -8.07 50.00 -38.83
Figure 5: (a) Global schematic diagram of MKG; (b) Local schematic diagram of MKG.
Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on publication year of paper. KG with earlier years applied similarity calculation algorithms to predict potential links, while KG with later years are used to verify how many sets of "Material Applications" in these predictions were reported in the following n years. The result shows in Figure 6. Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predication and the percentage of reported predication in next few year is displayed. The predictive capacity of the MKG is substantial, therefore, for each prediction, we only use the top 200 "material-application" pairs for validation. Looking at the results, as time progresses, the predicted materials are gradually verified. Using data from 2014, within nine years, 48.5% of the "material-application" predictions have been validated. This outcome demonstrates the effectiveness and reliability of the MKG, and it also lays the foundation for proposing new materials in some fields. The red, blue and green line are average percentage of reported link based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link predication, respectively. Network-based similarity has the highest accurate predication which indicates the effective of simulating the thinking of real material development using algorithms. The Figure 6 (b) is an example, visualizing the KG made from data before 2018, and the predicted "Material Application" reported within 5 years using network based algorithms.
Figure 6: Validation of the graph completion. (a) Percentage of reported predication after years. (b) Example of predicated material - application using data before 2018.
Conclusion and future work
In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts. This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets. Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text. This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted.
Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers. Our analysis confirms the effectiveness and credibility of the MKG. Furthermore, we have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding. This analysis has enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions. First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG. This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature. Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties. This enhancement significantly increases the granularity and practical utility of the KG. Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science. Lastly, the integration of our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset. This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.
Appendix
A. Validation for MKG
To validation of MKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, annotator can obtain 1000 entities and 500 relations. The report from annotator shows in Table 3. The labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous normalization of entities under these categories, using our expert dictionary. However, strict normalization can also lead to certain entity loss, which is difficult to avoid. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation. But the results are still satisfactory. "Name" and "Acronym" has also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. Fortunately, the impact of this kind of error is not significant and will not bring a significant influence, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
Table 3: Human evaluation metric on randomly selected triples.